# Task-Robust Pre-Training for Worst-Case Downstream Adaptation

### v1

@@ 4.1 NLP Scenario: Part-of-Speech Mask BERT
Task-Balancing MNLI QQP QNLI SST-2 CoLA STS-B MRPC RTE Avg.
None 84.6 71.2 90.5 93.5 52.2 85.8 88.9 66.4 79.6
@@ 4.1.2 Quantitative Result
Quantitative results are presented in Table 1. Our model obtains comparable results on GLUE tasks. PoS-BERT outperforms on half tasks by a substantial margin and obtains a 1.8% average score improvement over BERT. As for the most challenging training task, CoLA, which has the lowest accuracy on BERT, our method gets a 9.2% improvement which is a significant boost among downstream tasks. Benefiting from task-robust grouping, on QQP and RTE tasks, our method outperforms the original BERT by 5.7 F1-score and 4.3% accuracy, respectively. Our method also shows superiority on the challenging downstream task, MNLI, by achieving a 1.1% higher matched accuracy. As compensation for working better on the more challenging tasks, our method loses little correctness on some downstream tasks that already transfer well. On QNLI, SST-2, STS-B, and MRPC, our results are lower than that of the original BERT model by a margin of 1% accuracy, 1.9% accuracy, 1.6 spearson correlation, and 0.7 F1-score. The empirical result shows that, with our task-robust pre-training strategy, the downstream tasks perform on the whole, especially those tricky tasks. We expect future work to further improve these results by incorporating more sophisticated multi-task and grouping procedures.
@@ B.2 Analysis for the Minimax Pre-training Method
Appendix C Training details
@@
Hyperparam Part-of-Speech Mask BERT
Number of Layers 12
Hidden size 768
Attention heads 12
Attention heads size 64
Dropout 0.1
Warmup steps 10K
Peak Learning Rate 2e-4
Batch Size 256
Weight Decay 0.01
Max Steps 1000K
@@
Adam $\epsilon$ 1e-6
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.999
Gradient Clipping 0.0
Hyperparam GLUE Learning Rate 2e-5 Batch Size 32 Weight Decay 0.1 Learning Rate Decay Linear Warmup Ratio 0.06
@@
C.1 Part-of-Speech Mask BERT Training Setting
@@
C.2 Multi-Modal Mask MAE Training Setting

### v3

@@ 4.1 NLP Scenario: Part-of-Speech Mask BERT
Model Task-Balancing MNLI QQP QNLI SST-2 CoLA STS-B MRPC RTE Avg.
BERT - 84.6 71.2 90.5 93.5 52.2 85.8 88.9 66.4 79.6
PoS-BERT None 84.9 72.6 89.1 90.8 54.4 83.6 88.1 68.2 79.3
@@ 4.1.2 Quantitative Result
Quantitative results are presented in Table 1. Our model obtains comparable results on GLUE tasks. PoS-BERT with minimax task-balancing outperforms on half tasks by a substantial margin and obtains a 1.8% average score improvement over BERT. As for the most challenging training task, CoLA, which has the lowest accuracy on BERT, our method gets a 9.2% improvement which is a significant boost among downstream tasks. Benefiting from task-robust grouping, on QQP and RTE tasks, our method outperforms the original BERT by 5.7 F1-score and 4.3% accuracy, respectively. Our method also shows superiority on the challenging downstream task, MNLI, by achieving a 1.1% higher matched accuracy. As compensation for working better on the more challenging tasks, our method loses little correctness on some downstream tasks that already transfer well. On QNLI, SST-2, STS-B, and MRPC, our results are lower than that of the original BERT model by a margin of 1% accuracy, 1.9% accuracy, 1.6 spearson correlation, and 0.7 F1-score. The empirical result shows that, with our task-robust pre-training strategy, the downstream tasks perform on the whole, especially those tricky tasks. We expect future work to further improve these results by incorporating more sophisticated multi-task and grouping procedures.
@@ 6 Conclusion
7 Acknowledgments
C. Fang and Z. Lin were supported by National Key R&D Program of China (2022ZD0160301). C. Fang was also supported by the NSF China (No. 62376008) and Wudao Foundation. Z. Lin was also supported by the NSF China (No. 62276004), the major key project of PCL, China (No. PCL2021A12) an Qualcomm.
@@ B.2 Analysis for the Minimax Pre-training Method
Appendix C Example for Remark 5.2
Consider an extreme example where $\Theta=\mathbb{R}$, $\mathbb{E}_{z\sim P}\left[\ell_{1}(\theta,z)\right]=A(\theta-1)^{2}$ where $1<A<T$ and $\mathbb{E}_{z\sim P}\left[\ell_{t}(\theta,z)\right]=\theta^{2}$ for $t=2,\dots,T$. Then we have $\theta_{\text{max}}^{*}=\frac{\sqrt{A}}{\sqrt{A}-1}$ and $\theta_{\text{average}}^{*}=\frac{A}{A+T-1}$. It further implies that
$\ell_{\text{max}}$ $:=\max_{\lambda\in\Lambda}\mathbb{E}_{z\sim P}\left[\ell_{\lambda}\left(\theta_{\text{max}}^{*},z\right)\right]=\frac{A}{\left(\sqrt{A}-1\right)^{2}},$
$\ell_{\text{average}}$ $:=\max_{\lambda\in\Lambda}\mathbb{E}_{z\sim P}\left[\ell_{\lambda}\left(\theta_{\text{average}}^{*},z\right)\right]=\frac{A(T-1)^{2}}{(A+T-1)^{2}}.$
If $A=T-1$, we have $\frac{\ell_{\text{average}}}{\ell_{\text{max}}}\geq\frac{1}{4}\left(\sqrt{T-1}-1\right)^{2}$, which is as large as $O(T)$.
Appendix D Training details
@@
Hyperparam Part-of-Speech Mask BERT Number of Layers 12 Hidden size 768 Attention heads 12 Attention heads size 64 Dropout 0.1 Warmup steps 10K Peak Learning Rate 2e-4 Batch Size 256 Weight Decay 0.01 Max Steps 1000K Learning Rate Decay Linear Adam $\epsilon$ 1e-6 Adam $\beta_{1}$ 0.9 Adam $\beta_{2}$ 0.999 Gradient Clipping 0.0
Hyperparam GLUE
Learning Rate 2e-5
Batch Size 32
Weight Decay 0.1
@@
Warmup Ratio 0.06
@@
D.1 Part-of-Speech Mask BERT Training Setting
@@
D.2 Multi-Modal Mask MAE Training Setting